# Holographic entanglement negativity and thermodynamics in backreacted AdS black hole

### v1

@@ 6.2 High Temerature analysis
$\mathcal{E}\;\approx\;\lim_{B\to A^{c}}\frac{3}{8G_{N}}\left\{\frac{3R^{d-1}L^{d-2}}{2G_{N}}\left(\frac{4\pi T_{f}}{d}\right)^{d-2}\widetilde{S}(\rho,d)\right\},$ (6.6)
where $\widetilde{S}(\rho,d)$ is the dimensionless coefficient defined in eq.˜3.14 that increases monotonically with the backreaction parameter $\rho$. The important observation in this regimes are as follows
  • It exhibits the $T_{f}^{\,d-2}$ scaling shown in eq.˜6.6. This reflects the fact that, the dominant contribution to the relevant RT surfaces comes from regions of the bulk close to the horizon, so the negativity inherits the thermal (extensive) scaling of horizon-proximate geometric data.
[…]
The enhancement of $\mathcal{E}$ with $\rho$ at high temperature can be interpreted as follows: adding heavy-quark backreaction modifies the near-horizon geometry so that the bulk area-elements contributing to the negativity increase. Since high temperature already biases the system towards thermal (classical) correlations, the backreaction amplifies the same geometrical measure that feeds into the negativity, producing an overall larger value.
In the next section, we extend this analysis to the case of two disjoint subsystems, where the competition between thermal and quantum correlations manifests through a richer phase structure and provides further insight into the effect of backreaction on entanglement.
@@ 7 HEN for two disjoint subsystems
Let us denote two disjoint subsystems by $A_{1}$, $A_{2}$ and the region between them is by $A_{m}$ as follows
$x^{1}\in[-\frac{l_{1}}{2},\frac{l_{1}}{2}],\hskip 14.22636ptx^{i}\in[-\frac{L}{2},\frac{L}{2}]$ (7.1)
$x^{1}\in[-\frac{l_{2}}{2},\frac{l_{2}}{2}],\hskip 14.22636ptx^{i}\in[-\frac{L}{2},\frac{L}{2}]$ (7.2)
$x^{1}\in[-\frac{l_{m}}{2},\frac{l_{m}}{2}],\hskip 14.22636ptx^{i}\in[-\frac{L}{2},\frac{L}{2}]$ (7.3)
To calculate the HEN for two disjoint subsystems $A_{1}$ and $A_{2}$ we follow the the method outlined in [KumarBasak:2020viv]. Lets denotes the turning points by $z_{t_{i}}=z_{t_{1}},z_{t_{2}},z_{t_{m}}$ of the RT surfaces anchored from the boundary of $A_{1}$, $A_{2}$ and $A_{m}$ respectively. The HEN for this disjoint case is given by
$\mathcal{E}=\frac{3}{16G_{N}}(\mathcal{A}_{A_{1}\cup A_{m}}+\mathcal{A}_{A_{m}\cup A_{2}}-\mathcal{A}_{A_{1}\cup A_{m}\cup A_{2}}-\mathcal{A}_{A_{m}})$ (7.4)
Now we will only focus on two regimes as done in previous section.
@@ 7.1 Low temperature limit
Following the construction introduced in eq.˜3.11 and eq.˜7.4. At low effective temperature ($T_{f}l_{j}\ll 1$), one obtains the analytic expression for HEN as follow
$\begin{split}\mathcal{E}=\frac{3R^{d-1}L^{d-2}}{16G_{N}}S_{0}\Biggl[&(l_{1}+l_{m})^{2-d}+(l_{2}+l_{m})^{2-d}-(l_{1}+l_{2}+l_{m})^{2-d}-(l_{m})^{2-d}\\ &\hskip 8.5359pt-\,2S_{0}S_{1}(1-\rho)\,l_{1}l_{2}\left(\frac{4\pi T_{f}}{d}\right)^{\!d}+\mathcal{O}\!\left((T_{f}l_{j})^{2d-2}\right)\Biggr],\end{split}$ (7.5)
where $S_{0}$ and $S_{1}$ are $d$-dependent constants previously defined.
It is instructive to verify that in the limit $l_{m}\to 0$, the above expression correctly reproduces the result for the adjacent interval configuration obtained in eq.˜5.4. Indeed, setting $l_{m}=0$ in eq.˜7.5 yields eq.˜5.4 which is precisely the low-temperature adjacent HEN result.
From eq.˜7.5 one can observe that
@@
  • The backreaction enters explicitly through the correction term proportional to $(1-\rho)\,T_{f}^{\,d}$. Similar to the adjacent case, this term suppresses the negativity, indicating that the presence of heavy quarks diminishes quantum correlations at low temperature by enhancing thermal effects.
  • The separation width $l_{m}$ plays a nontrivial role: as $l_{m}$ increases, the negativity decreases due to reduced overlap of the entanglement wedges, consistent with the expectation that large separations weaken correlations.
In summary, the low-temperature behavior of HEN for disjoint intervals captures both the universal short-distance scaling and the suppressive effect of backreaction. In the next section, we will analyze the high-temperature regime, where the interplay between thermal contributions and $\rho$-dependent corrections leads to qualitatively different behavior.
@@ 7.2 High temperature limit
We now turn to the analysis of HEN for two disjoint subsystems in the high-temperature regime under $l_{m}\ll l$ condition. Following the procedure outlined in [KumarBasak:2020viv], and employing the general expressions eqs.˜3.13 and 7.4, one obtains
$\mathcal{E}=\frac{3L^{d-2}R^{d-1}}{8G_{N}}\left(\frac{4\pi T_{f}}{d}\right)^{\!d}\Bigl[\widetilde{S}_{1m}(\rho,d)+\widetilde{S}_{2m}(\rho,d)-\widetilde{S}_{1m2}(\rho,d)\Bigr]-\frac{3}{4}S_{A_{m}}^{low},$ (7.6)
where the coefficients $\widetilde{S}(\rho,d)$ are defined in eq.˜3.14 and $S_{A_{m}}^{\mathrm{low}}$ is given by eq.˜4.29.
As a nontrivial check of consistency, we can take the limit $l_{m}\to 0$ in (7.6), which yields
$\mathcal{E}=\frac{3L^{d-2}R^{d-1}}{8G_{N}}\left(\frac{4\pi T_{f}}{d}\right)^{\!d}\Bigl[\widetilde{S}_{1}(\rho,d)+\widetilde{S}_{2}(\rho,d)-\widetilde{S}_{12}(\rho,d)\Bigr],$ (7.7)
precisely matching the adjacent interval result given in eq.˜5.5. This provides a strong consistency check for the validity of the prescription. The following are the main observations for this case:
  • The negativity scales as $T_{f}^{\,d}$, in contrast to the $T_{f}^{\,d-2}$ scaling observed for the adjacent case. This enhanced scaling originates from the explicit dependence of the disjoint configuration on the separation width $l_{m}$.
  • The quantities $\widetilde{S}_{1m}(\rho,d)$, $\widetilde{S}_{2m}(\rho,d)$, $\widetilde{S}_{1m2}(\rho,d)$, and $\widetilde{S}_{m}(\rho,d)$ capture the geometric and backreaction-dependent contributions. Each of these functions increases monotonically with $\rho$, implying that the negativity is enhanced as the quark density grows. This behavior is consistent with the previous cases at high temperature but is in sharp contrast with the suppression observed in the low-temperature regime.
  • The separation width $l_{m}$ introduces an additional control parameter: increasing $l_{m}$ suppresses the negativity, reflecting the weakening of correlations between spatially separated subsystems. The balance between this geometric suppression and the $\rho$-induced enhancement determines the overall behavior of the negativity in this regime.
The analysis of (7.5) and (7.6) reveals the behavior of HEN for disjoint subsystems across thermal regimes. At low temperature, the leading $\rho$-dependent correction appears at order $T_{f}^{\,d}$ and acts to suppress the negativity, consistent with the intuition that heavy quark backreaction enhances thermal correlations at the cost of quantum correlations. In contrast, at high temperature the negativity scales as $T_{f}^{\,d}$ with coefficients $\widetilde{S}(\rho,d)$ that grow monotonically with $\rho$, resulting in an enhancement of quantum correlations in the thermal phase. The separation width $l_{m}$ further modulates this behavior, reducing correlations as the subsystems are pulled apart, but never altering the qualitative difference between the two thermal regimes. This comparative picture establishes that the influence of backreaction is highly regime-dependent: it diminishes entanglement at low temperature while enhancing it at high temperature.
8 Critical separation and entanglement–disentanglement transition
@@
In this section we will compute the critical separation between two subregion in disjoint configuration. In [Chakrabortty:2020ptb], the critical separation between two subsystems is studied using holographic mutual information. The critical separation is defined as the value of the intermediate width at which the HEN vanishes. The HEN for disjoint subsystems from eq.˜7.4 can be rewritten as,
@@
At critical width $l_{m}=l_{c}$ we have,
@@
We focus on the physically relevant regime $l_{c}\ll l_{1},l_{2}$, where the entropies associated with the larger subsystems probe the high-temperature limit, while the entropy corresponding to the separation region remains in the low-temperature regime. Accordingly,
@@
From the low-temperature expansion of the holographic entanglement entropy eq.˜3.11, only the leading non-analytic term $S_{0}\,l^{2-d}$ survives for the narrow separation region $l=l_{c}$, since all analytic thermal corrections are parametrically suppressed for $l_{c}T_{f}\ll 1$. Moreover, the UV-divergent term cancels identically in the negativity combination. Therefore, to leading order,
$S^{\text{low}}(l_{c})=\frac{R^{d-1}L^{d-2}}{2G_{N}}\Big[S_{0}\,l_{c}^{2-d}+\mathcal{O}(l_{c}T_{f}^{d-1})\Big].$ (8.4)
The high-temperature HEE for a strip of width $l$ is given by eq.˜3.13 and can be rewritten as,
$S^{\text{high}}(l)=\frac{R^{d-1}L^{d-2}}{2G_{N}}\Big[\alpha_{d}\,T_{f}^{d-2}\,l+\beta_{d}(\rho)\,T_{f}^{d-2}\Big],$ (8.5)
with the effective coefficients
$\alpha_{d}=\frac{1}{2}\left(\frac{4\pi}{d}\right)^{d-1},$ (8.6)
$\beta_{d}(\rho)=\widetilde{S}(\rho,d)\left(\frac{4\pi}{d}\right)^{d-2}.$ (8.7)
The coefficient $\alpha_{d}$ controls the universal extensive thermal contribution, while $\beta_{d}(\rho)$ captures the finite backreaction-dependent correction arising from the near-horizon geometry. Using (8.5), the combination of high-temperature entropies appearing in (8.1) simplifies to
$\begin{split}&S^{\text{high}}(l_{1}+l_{c})+S^{\text{high}}(l_{2}+l_{c})-S^{\text{high}}(l_{1}+l_{2}+l_{c})\\ &\qquad=\frac{R^{d-1}L^{d-2}}{2G_{N}}\Big[\alpha_{d}\,T_{f}^{d-2}\,l_{c}+\beta_{d}(\rho)\,T_{f}^{d-2}\Big],\end{split}$ (8.8)
where all dependence on $l_{1}$ and $l_{2}$ cancels exactly. Imposing the condition (8.2) and combining (8.4) with (8.8), we obtain the algebraic equation
$\alpha_{d}\,T_{f}^{d-2}\,l_{c}+\beta_{d}(\rho)\,T_{f}^{d-2}=S_{0}\,l_{c}^{2-d}\left[1+\mathcal{O}(l_{c}T_{f})\right].$ (8.9)
In the regime $l_{c}T_{f}\ll 1$, the power-law term dominates, yielding the leading-order solution
$l^{(0)}_{c}=\left(\frac{S_{0}}{\beta_{d}(\rho)\,T_{f}^{d-2}}\right)^{\!\frac{1}{d-2}}.$ (8.10)
Since $\beta_{d}(\rho)$ increases monotonically with the backreaction parameter $\rho$, the critical separation decreases as $\rho$ increases. This indicates that backreaction suppresses long-range distillable entanglement, leading to an earlier onset of the entanglement–disentanglement transition. This behavior is in sharp contrast with the enhancement observed for the EWCS and reflected entropy at low temperature, underscoring the distinct physical information captured by holographic entanglement negativity.
In [Chakrabortty:2020ptb], the critical separation between two disjoint strips was analyzed using holographic mutual information (HMI), where the transition point corresponds to a change in the connectivity of the entanglement wedge. The resulting critical distance reflects the persistence of total correlations, including both quantum and classical contributions. In contrast, the critical separation $l_{c}$ obtained from HEN characterizes the vanishing of distillable quantum entanglement between the subsystems. As a consequence, the negativity-based critical distance is generically smaller than that obtained from HMI. This hierarchy is physically expected: while classical correlations can survive at larger separations and temperatures, quantum entanglement is more fragile and is suppressed more efficiently by thermal and backreaction effects. In the present backreacted AdS black hole background, increasing the quark density parameter $\rho$ enhances the thermal entropy and accelerates the disentanglement of quantum correlations, leading to a monotonic decrease of $l_{c}$ from negativity. On the other hand, the HMI-based critical distance exhibits a comparatively weaker suppression with $\rho$, consistent with its sensitivity to classical correlations encoded in the geometry of the entanglement wedge. The comparison thus highlights the complementary roles of HMI and HEN: while both diagnose correlation phase transitions, negativity provides a stricter and more refined probe of quantum entanglement in strongly coupled plasmas.

### v2

@@ 6.2 High temperature analysis
$\mathcal{E}=\frac{3\,L^{d-2}R^{d-1}}{4G_{N}^{d+1}}\left[\frac{1}{(d-2)\epsilon^{d-2}}+\biggl(\frac{4\pi T_{f}}{d}\biggr)^{\!d-2}\,\widetilde{S}(\rho,d)\right],$ (6.6)
where $\widetilde{S}(\rho,d)$ is defined in eq.˜3.14 and it increases monotonically with the backreaction parameter $\rho$. The above expression can also be written in the following form
$\mathcal{E}=\frac{3}{2}\left(S_{A}-S_{A}^{th}\right).$ (6.7)
Consistent with the previous case, this behavior indicates the elimination of extensive thermal components from the negativity. The important observation in this regimes are as follows:
  • We note that the leading order expression for the HEN in eq.˜6.6 depends only on the area of the entangling surface $L^{d-2}$. This area law scaling for the mixed state negativity is consistent with the findings reported in various systems like finite-temperature quantum spin models and two-dimensional harmonic lattices [22, 75].
  • The HEN increases with the backreaction parameter $\rho$ through the monotonic growth of the coefficient $\widetilde{S}(\rho,d)$. Apart from the ultraviolet divergent term, the leading finite contribution exhibits a temperature dependence scaling as $T_{f}^{\,d-2}$.
  • Similar to the adjacent configuration, the enhancement of entanglement negativity with increasing backreaction reflects the presence of additional degrees of freedom in the system. However, unlike the HEE which scales with the subsystem volume and becomes insensitive to quantum correlations at high temperature, the entanglement negativity remains a reliable probe of quantum correlations in this regime.
Thus the bipartite analysis reinforces the distinctive role of HEN in isolating quantum correlations in mixed states. Even in regimes where thermal effects and backreaction are significant, the negativity continues to capture entanglement features that remain sensitive to the underlying degrees of freedom thereby providing a complementary and more faithful diagnostic than other entanglement measures at high temperature.
@@ 7 HEN for two Disjoint Subsystems
We now analyze the HEN for the case of mixed state configuration of two disjoint subsystems. This configuration is described by two long rectangular strips $A_{1}$ and $A_{2}$ separated by an another strip $A_{m}$ as follows
$x\equiv x^{1}\in[-\frac{l_{j}}{2},\frac{l_{j}}{2}],\quad x^{i}\in[-\frac{L}{2},\frac{L}{2}];\quad i=2,3,\dots,(d-1),\quad j=1,2,m,$ (7.1)
where $L\gg l_{1},l_{2},l_{m}$ in order to neglect the corner effects. To determine the HEN for the two disjoint subsystems $A_{1}$ and $A_{2}$, we follow the the method outlined in [54]. The HEN for the disjoint case can be expressed as
$\mathcal{E}=\frac{3}{16G_{N}^{d+1}}(\mathcal{A}_{A_{1}\cup A_{m}}+\mathcal{A}_{A_{m}\cup A_{2}}-\mathcal{A}_{A_{1}\cup A_{m}\cup A_{2}}-\mathcal{A}_{A_{m}}),$ (7.2)
where $\mathcal{A}_{A_{i}\cup A_{j}}$ and $\mathcal{A}_{A_{i}\cup A_{j}\cup A_{k}}$ denotes the area of the RT surfaces for the subsystems $A_{i}\cup A_{j}$ and $A_{i}\cup A_{j}\cup A_{k},$ respectively with $i,j,k=1,2,m$. Now we will focus on two regimes corresponding to low and high effective temperature limit as done in previous section.
@@ 7.1 Low temperature limit
In the low effective temperature regime ($T_{f}l_{j}\ll 1$), we obtain the analytic expression for HEN of two disjoint subsystems by using eq.˜3.11 in eq.˜7.2 as follows
$\begin{split}\mathcal{E}=\frac{3R^{d-1}L^{d-2}}{16G_{N}^{d+1}}S_{0}\Biggl[&(l_{1}+l_{m})^{2-d}+(l_{2}+l_{m})^{2-d}-(l_{1}+l_{2}+l_{m})^{2-d}-(l_{m})^{2-d}\\ &\hskip 8.5359pt-\,2S_{1}(1-\rho)\,l_{1}l_{2}\left(\frac{4\pi T_{f}}{d}\right)^{\!d}+\mathcal{O}\!\left((T_{f}l_{j})^{2d-2}\right)\Biggr],\end{split}$ (7.3)
where $S_{0}$ and $S_{1}$ are $d$-dependent constants as defined earlier in eq.˜3.10. It is instructive to verify that in the adjacent limit $l_{m}\to 0$, the above expression reproduces the HEN for adjacent subsystems obtained in eq.˜5.4. Further, the above result matches with the corresponding HEN for disjoint subsystems in the absence of the backreaction $(\rho\to 0)$ [54]. From eq.˜7.3 one can observe that:
  • In contrast to the adjacent and bipartite configurations, the ultraviolet divergent contributions cancel completely in this case, leaving the HEN finite. Consequently, the negativity scales with the area of the subsystem and obeys an area law.
@@
  • The backreaction enters explicitly through the correction term proportional to $(1-\rho)\,T_{f}^{\,d}$. Similar to the adjacent case, this term increases the negativity with increasing backreaction. This indicates that the presence of heavy quarks or extra degrees of freedoms enhances the quantum correlations at low temperature.
  • The separation width $l_{m}$ plays a nontrivial role as $l_{m}$ increases, the negativity decreases due to reduced overlap of the entanglement wedges which is consistent with the expectation that large separations weaken correlations.
In summary, the low temperature behavior of HEN for disjoint subsystems captures both the universal short distance scaling and the effect of backreaction. In the next section, we will analyze the high temperature regime where the interplay between thermal contributions and $\rho$-dependent corrections leads to qualitatively different behavior.
@@ 7.2 High temperature limit
We now turn to the analysis of HEN for mixed state configuration of two disjoint subsystems in the high effective temperature regime under $l_{m}\ll l_{1,2}$ condition. In this case we utilize the high temperature result eq.˜3.13 for all area functions appearing in eq.˜7.2 except $\mathcal{A}_{m}$, for which the low temperature result eq.˜3.11 is applicable. Now employing the general expressions from eqs.˜3.11 and 3.13 in eq.˜7.2 lead to the following expression for HEN as
$\mathcal{E}=\frac{3L^{d-2}R^{d-1}}{16G_{N}^{d+1}}\Biggl[2\biggl(\frac{4\pi T_{f}}{d}\biggr)^{d-2}\,\widetilde{S}(\rho,d)$ $+l_{m}\biggl(\frac{4\pi T_{f}}{d}\biggr)^{d-1}(1-S_{0}S_{2}\rho)$ (7.4)
$-S_{0}S_{1}(1-\rho)\,l_{m}^{2}\left(\frac{4\pi T_{f}}{d}\right)^{d}-S_{0}\,l_{m}^{2-d}\Biggl],$
The above expression is independent of $l_{1}$ and $l_{2}$ and contain only the width $l_{m}$ of the intermediate region $A_{m}$. The following are the main observations for this case:
  • The term proportional to volume contributions from subsystems $A_{1}$ and $A_{2}$ get completely subtracted out as they are in large effective temperature regime.
  • The first finite term scale as $T_{f}^{\,d-2}$ and encoding the nontrivial dependence on the backreaction and spacetime dimension through $\widetilde{S}(\rho,d)$. This term captures quantum correlations that persist beyond the leading thermal behavior and suggest an increase in correlation with backreaction.
  • Next term is linear in the separation width $l_{m}$ and scales as $T_{f}^{\,d-1}$, indicating that at high temperature the entanglement negativity increases with separation between subsystems. The factor $(1-S_{0}S_{2}\rho)$ shows that increasing backreaction suppresses this contribution.
  • The negative contributions proportional to $l_{m}^{2}$ and $l_{m}^{2-d}$ reflect the competing effects of thermal and geometric suppression at large separation. In particular, the $(1-\rho)\,l_{m}^{2}T_{f}^{\,d}$ term implies that increasing backreaction further increase the negativity.
Overall, the high-temperature entanglement negativity for disjoint subsystems exhibits a rich interplay between thermal effects, geometric separation and matter backreaction. While contributions proportional to $\widetilde{S}(\rho,d)$ encode persistent quantum correlations that are enhanced by backreaction, the separation-dependent terms reveal competing mechanisms that either decrease or amplify the negativity depending on the parametric regime. In the following section, we analyze how these competing contributions determine a critical separation beyond which the entanglement negativity of the disjoint configuration vanishes.
8 Critical Separation and Entanglement Disentanglement Transition
@@
In this section we will compute the critical separation between two subregions in disjoint configuration. In [14], the effect of backreaction on critical separation between two subsystems is studied using holographic mutual information and it was observed that the critical separation increases with backreaction. Since the HEN is also an entanglement measure for quantum correlations, we can use it to detect the critical separation or phase transition from entangled to disentangled phase. The critical separation is defined as the value of the intermediate width at which the HEN vanishes. The HEN for disjoint subsystems from eq.˜7.2 can be rewritten as
@@
At critical width $l_{m}=l_{c}$, we have
@@
We focus on the physically relevant regime $l_{c}\ll l_{1},l_{2}$, where the entropies associated with the larger subsystems probe the high temperature limit while the entropy corresponding to the separating region remains in the low temperature regime. Accordingly,
@@
From the high and low temperature expansions of the HEE and imposing eq.˜8.2 we obtain,
$\begin{split}2\widetilde{S}(\rho,d)\left(\frac{4\pi}{d}\right)^{d-2}(T_{f}l_{c})^{d-2}+(1-S_{0}S_{2}\rho)&\left(\frac{4\pi}{d}\right)^{d-1}(T_{f}l_{c})^{d-1}\\ &+S_{0}S_{1}(1-\rho)\left(\frac{4\pi}{d}\right)^{d-2}(T_{f}l_{c})^{d}=S_{0}.\end{split}$ (8.4)
We now focus on the regime $T_{f}l_{c}\ll 1$ which corresponds to small separations compared to the thermal scale. In this limit, the $(T_{f}l_{c})^{d-2}$ term dominates over the $(T_{f}l_{c})^{d-1}$ and $(T_{f}l_{c})^{d}$ terms, and eq.˜8.4 reduces at leading order to
$2\widetilde{S}(\rho,d)\left(\frac{4\pi}{d}\right)^{d-2}(T_{f}l_{c})^{d-2}\simeq S_{0}.$ (8.5)
Solving for the critical separation, we obtain
$l^{(0)}_{c}\simeq\frac{1}{T_{f}}\left[\frac{S_{0}}{2\widetilde{S}(\rho,d)}\left(\frac{d}{4\pi}\right)^{d-2}\right]^{\frac{1}{d-2}}.$ (8.6)
Since $\widetilde{S}(\rho,d)$ is a monotonically increasing function of the backreaction parameter $\rho$ and $S_{0}$ is negative constant, eq.˜8.6 as noted in [14], $\widetilde{S}(0,4)=-0.33$ and $\widetilde{S}(1,4)=-0.024$ therefore it immediately implies that in the range of backreaction $\rho=0$ to $\rho=1$ we have
$\frac{\partial l_{c}}{\partial\rho}>0.$ (8.7)
However, to see this increasing behavior for full range of $\rho$ we need to include all terms. Thus, in the regime $T_{f}l_{c}\ll 1$, the critical separation extracted from entanglement negativity increases with increasing backreaction up to a certain value of $\rho$ after a particular value of backreaction critical separation becomes imaginary which requires further computation with subleading corrections in eq.˜8.4.
This behavior is similar with the critical separation obtained from mutual information in [14]. The critical separation characterizes the spatial range over which distillable quantum correlations persist. Matter backreaction enhances thermal mixing and accelerates the decay of quantum correlations with distance, causing the negativity to vanish. While mutual information diagnoses the persistence of total correlations, entanglement negativity provides a sharper probe of the decay of quantum entanglement in strongly coupled holographic systems with matter backreaction.